# Uncertainty in Language Models: Assessment through Rank-Calibration

### v1

@@ Abstract
Language Models (LMs) have shown promising performance in natural language generation. However, as LMs often generate incorrect or hallucinated responses, it is crucial to correctly quantify their uncertainty in responding to given inputs. In addition to verbalized confidence elicited via prompting, many uncertainty measures (e.g., semantic entropy and affinity-graph-based measures) have been proposed. However, these measures can differ greatly, and it is unclear how to compare them, partly because they take values over different ranges (e.g., $[0,\infty)$ or $[0,1]$). In this work, we address this issue by developing a novel and practical framework, termed Rank-Calibration, to assess uncertainty and confidence measures for LMs. Our key tenet is that higher uncertainty (or lower confidence) should imply lower generation quality, on average. Rank-calibration quantifies deviations from this ideal relationship in a principled manner, without requiring ad hoc binary thresholding of the correctness score (e.g., ROUGE or METEOR). The broad applicability and the granular interpretability of our methods are demonstrated empirically.
@@ 1 Introduction
Language Models (LMs), especially Large Language Models (LLMs), have shown promising performance in Natural Language Generation (NLG). These models, fitted on huge text corpora, can produce responses resembling those of humans Touvron et al. (2023b); OpenAI (2023). However, since LMs often generate wrong or hallucinated responses (Weidinger et al., 2021; Xiao and Wang, 2021), it is crucial to correctly quantify their level of uncertainty in responding to particular inputs.
Refer to captionRefer to caption
@@ 2 Correctness and Uncertainty for LMs
We work with a deterministic correctness function $A:\mathcal{V}^{\star}\times\mathcal{V}^{\star}\to\mathbb{R}$ mapping each pair $(\boldsymbol{x};\widehat{\boldsymbol{y}})$ to a correctness value $A(\boldsymbol{x};\widehat{\boldsymbol{y}})$. In practice, correctness is often not a binary variable in NLG tasks and can be assessed in at least two different ways.
@@ 3 Limitations of Existing Assessments
The uncertainty measures examined here include the negative log-likelihood $U_{\rm NLL}$, the semantic entropy $U_{\rm SE}$ Kuhn et al. (2023), the affinity-graph-based measures $U_{\rm EigV}$, $U_{\rm Ecc}$, and $U_{\rm Deg}$ (Lin et al., 2023), with affinity determined by the NLI model He et al. (2021), and the verbalized confidence $C_{\rm Verb}$ (Xiong et al., 2024); see definitions in Appendix B. These include both white box and grey box measures, [^3] The grey-box oracle refers to the access to model logits, which is partly feasible for commercial LMs, while the black-box oracle only relies on generated outputs. as well as a diversity of prompt strategies. We use the Rouge-L score as the correctness function $A$. We follow a common assessment pipeline (Kuhn et al., 2023; Lin et al., 2023; Xiong et al., 2024), as depicted in Fig. 2. The assessment metrics are detailed in Appendix C.
@@
Most existing assesment metrics (e.g., AUROC, AUPRC, ECE, etc) are rooted in classification and require binary labels (i.e., $A\in\{\textit{True}\text{ or }\textit{False}\}$). Consequently, an ad hoc threshold $\tau\in\mathbb{R}$ is often introduced to map continuous correctness values to binary labels, i.e., $\bar{A}_{\tau}(\boldsymbol{x};\widehat{\boldsymbol{y}}):=\mathds{1}[A(\boldsymbol{x};\widehat{\boldsymbol{y}})\geq\tau]$ Lin et al. (2023); Kuhn et al. (2023). Thus, the response is viewed as correct if the correctness value $A(\boldsymbol{x};\widehat{\boldsymbol{y}})$ is at least $\tau$, and incorrect otherwise.
@@
The second limitation of existing assessments is rooted in the diverse output ranges of the uncertainty or confidence measures. As shown in Fig. 3 [bottom], the output ranges of different uncertainty measures vary significantly. For example, the values of $U_{\rm SE}$ can be higher than $100$ while the values of $U_{\rm Ecc}$ and $U_{\rm Deg}$ are small by definition. This diversity of output ranges prevents the direct use of calibration-based metrics such as the ECE, which takes variables with inputs in $[0,1]$.
@@ 4.1 Rank-Calibration & RCE
Since $U$ is a random variable depending on $(\boldsymbol{x};\widehat{\boldsymbol{y}})$, $\mathrm{reg}(U)$ is also random. If $\mathrm{reg}(\cdot)$ is monotone decreasing, then $U\leq u$ implies $\mathrm{reg}(U)\geq\mathrm{reg}(u)$. Thus, for any value $u$ in the range of $U$,
$\vspace{-1mm}\mathbb{P}(U\leq u)=\mathbb{P}(\mathrm{reg}(U)\geq\mathrm{reg}(u)).$ (7)
@@
We say that an uncertainty measure $U$ is rank-calibrated if (7) holds for any $u$ in $U$’s range: on average, lower uncertainty implies higher generative quality.
Rank-calibration is related to, yet distinct from, the usual notion of calibration in the classification context (Lichtenstein et al., 1977; Guo et al., 2017). We defer the detailed discussion to Sec. 4.2.
@@ 4.2 Comparison with Classical Calibration
Theorem 1 implies that, for a binary correctness function, one can construct a calibrated confidence measure from an uncertainty measure with monotone transformations if and only if the uncertainty measure is rank-calibrated. However, RCE and ECE gauge different quantities: ECE captures the absolute difference between predicted and true probabilities, while RCE reflects the deviation from a monotonic correspondence between uncertainty and the expected correctness. These two notions are generally not directly comparable.
For example, consider the special case where a continuous-valued confidence measure $C$ is completely uninformative and the regressed correctness $\overline{\mathrm{reg}}:c\mapsto\mathbb{E}[A\mid C=c]$ is a constant for all confidence level $c$. Then, the RCE defined in (8) reports a large value of $1/2$, reflecting its poor indicativeness. However, the ECE can be large or small depending on the averaged distance between $C$’s output and $\overline{\mathrm{reg}}$. More generally, we find no relation in the results of ECE and RCE through the following result, proved in Appendix D.
@@ 4.3 Empirical RCE & Indication Diagram
To address this, we adopt a piecewise constant regression or binning strategy, as in non-parametric statistics Tsybakov (2009). First, we group the uncertainty values $\{u_{i}\}_{i=1}^{n}$ into $B$ equal-mass intervals, each containing $\lceil n/B\rceil$—or, when needed, $\lfloor n/B\rfloor$—elements. The boundaries of the $b$-th ($1\leq b\leq B$) bin are the $(b-1)/B$-th and $b/B$-th quantiles of $(u_{i})_{i=1}^{n}$. Let $\mathcal{I}_{b}\subseteq\{1,\dots,n\}$ be the set of indices of the datapoints whose uncertainty values fall into the $b$-th bin. Then, the expected correctness level over the $b$-th bin can estimated as
$\vspace{-2mm}{\rm crc}_{b}:=\frac{1}{|\mathcal{I}_{b}|}\sum_{i\in\mathcal{I}_{b}}a_{i},$
@@
$\vspace{-2mm}{\rm uct}_{b}=\frac{1}{|\mathcal{I}_{b}|}\sum_{i\in\mathcal{I}_{b}}u_{i}.$
As ${\rm crc}_{b}$ and ${\rm uct}_{b}$ estimate the per-bin averages of $\mathrm{reg}(U)$ and $U$, for each $b$. We estimate $\mathbb{P}(U\leq u_{i})$ and $\mathbb{P}(\mathrm{reg}(U)\geq\mathrm{reg}(u_{i}))$ for $i\in\mathcal{I}_{b}$ as follows:
^$\widehat{\mathbb{P}}(\mathrm{reg}(U)\geq\mathrm{reg}(u_{i})):=\frac{1}{B-1}\sum_{b^{\prime}\neq b}\mathds{1}[{\rm crc}_{b}\geq{\rm crc}_{b^{\prime}}],\vspace{-1mm}$
^$\widehat{\mathbb{P}}(U\leq u_{i}):=\frac{1}{B-1}\sum_{b^{\prime}\neq b}\mathds{1}[{\rm uct}_{b}\leq{\rm uct}_{b^{\prime}}].$
@@ 5.1 Experiment Setup
The uncertainty/confidence measures to be assessed are the same as in Sec. 3, (i.e., $U_{\rm NLL}$, $U_{\rm SE}$, $U_{\rm Ecc}$, $U_{\rm Deg}$, $U_{\rm EigV}$, and $C_{\rm Verb}$). We first illustrate that our proposed assessment has broad applicability and granular interpretability. Furthermore, we qualitatively show that uncertainty measures with lower RCE values reliably indicate correctness. Finally, we study robustness by empirically checking the impact of temperature and correctness functions on RCE (Demšar, 2006). More results for different configurations are in Table 2.
@@ 5.5 Post-hoc Recalibration
Recalibrating uncertainty/confidence measures with poor rank-calibration can be of interest; for ECE, this is sometimes known as Mincer-Zamowitz regression Mincer and Zarnowitz (1969). As discussed in Sec. 4.2, an ECE-calibrated measure is also RCE-calibrated. However, RCE is invariant to monotone transformations, which means that approaches like Platt scaling (Platt, 1999) and isotonic regression (Zadrozny and Elkan, 2002) will not improve rank-calibration. Therefore, we suggest using histogram binning (or, piecewise constant regression), which includes non-monotone transforms (Zadrozny and Elkan, 2001). See the related experimental results in Appendix F.2.
@@ 5.6 Robustness Analysis
We conduct ablation studies to analyze the robustness of our assessment to key hyperparameters, including temperatures and correctness scores. We further propose a method to make robust comparisons between uncertainty measures via the Critical Difference (CD) Diagram  (Demšar, 2006). Detailed information and results are in Appendix F.4.
@@ 6 Conclusion
Appendix A More Related Work
@@
The quantification of uncertainties in model outputs in supervised learning has a long history (e.g., Lichtenstein et al., 1977, etc). Overparametrized models such as neural networks pose unique challenges to estimating uncertainty and enhancing model calibration(Guo et al., 2017; Papadopoulos et al., 2001; Riquelme et al., 2018). Various approaches have been introduced to mimic Bayesian inference (Gal and Ghahramani, 2016), to utilize simple deep ensembles (Lakshminarayanan et al., 2017; Jain et al., 2020), and to identify training samples that are out-of-distribution (Liang et al., 2018; Papernot and McDaniel, 2018). Nonetheless, it is not clear how to adapt these strategies to language modeling, where the outputs can be text with complex structure.
@@
To gauge the uncertainty level associated with the outputs of LMs, Kuhn et al. (2023) introduce the concept of semantic entropy, which integrates linguistic consistencies stemming from shared meanings. In a similar vein, Kadavath et al. (2022); Lin et al. (2022); Xiong et al. (2024) encourage LMs to analyze their own responses and come up with a “probability” that a response is correct. In related work, Manakul et al. (2023) uses sampling to identify instances of fabricated information. Recently, Tian et al. (2023) explore methods for deriving confidence measures for reinforcement-learning-trained LMs. Lin et al. (2023) draw a distinction between estimating uncertainty and confidence for LMs. Similarly, Chen and Mueller (2023) introduce a method for detecting bad and speculative responses from a pre-trained LM with a confidence score. Tanneru et al. (2023) propose two novel measures to quantify the uncertainty of LM-generated explanations. Although considerable research focuses on developing uncertainty and confidence measures for LMs, the evaluation of their effectiveness is less studied.

### v2

@@ Abstract
Language Models (LMs) have shown promising performance in natural language generation. However, as LMs often generate incorrect or hallucinated responses, it is crucial to correctly quantify their uncertainty in responding to given inputs. In addition to verbalized confidence elicited via prompting, many uncertainty measures (e.g., semantic entropy and affinity-graph-based measures) have been proposed. However, these measures can differ greatly, and it is unclear how to compare them, partly because they take values over different ranges (e.g., $[0,\infty)$ or $[0,1]$). In this work, we address this issue by developing a novel and practical framework, termed Rank-Calibration, to assess uncertainty and confidence measures for LMs. Our key tenet is that higher uncertainty (or lower confidence) should imply lower generation quality, on average. Rank-calibration quantifies deviations from this ideal relationship in a principled manner, without requiring ad hoc binary thresholding of the correctness score (e.g., ROUGE or METEOR). The broad applicability and the granular interpretability of our methods are demonstrated empirically. The code to replicate our experiments is here.
@@ 1 Introduction
Language Models (LMs), especially Large Language Models (LLMs), have shown promising performance in Natural Language Generation (NLG). These models, fitted on huge text corpora, can produce responses resembling those of humans Touvron et al. (2023b); OpenAI (2023). However, since LMs often generate wrong or hallucinated responses (Weidinger et al., 2021; Xiao and Wang, 2021; Huang et al., 2024), it is crucial to correctly quantify their level of uncertainty in responding to particular inputs.
@@ 2 Correctness and Uncertainty for LMs
We work with a deterministic correctness function $A:\mathcal{V}^{\star}\times\mathcal{V}^{\star}\to\mathbb{R}$ mapping each pair $(\boldsymbol{x};\widehat{\boldsymbol{y}})$ to a correctness value $A(\boldsymbol{x};\widehat{\boldsymbol{y}})$. In practice, correctness is often not a binary variable in NLG tasks and can be assessed in at least two different ways. For the reader’s convenience, the concepts and notations used in the paper are summarized in Table 1.
@@
The closest works have been discussed in Sec. 1 and 2, and more related works are reviewed in Appendix A.
@@ 3 Limitations of Existing Assessments
The uncertainty measures examined here include the negative log-likelihood $U_{\rm NLL}$, the semantic entropy $U_{\rm SE}$ Kuhn et al. (2023), the affinity-graph-based measures $U_{\rm EigV}$, $U_{\rm Ecc}$, and $U_{\rm Deg}$ (Lin et al., 2023), with the affinity determined by the NLI model He et al. (2021), and the verbalized confidence $C_{\rm Verb}$ (Xiong et al., 2024); see the definitions in Appendix B. These include both white box and grey box measures, [^3] The grey-box oracle refers to the access to model logits, which is partly feasible for commercial LMs, while the black-box oracle only relies on generated outputs. as well as a diversity of prompt strategies. We use the Rouge-L score as the correctness function $A$. We follow a common assessment pipeline (Kuhn et al., 2023; Lin et al., 2023; Xiong et al., 2024), as depicted in Fig. 2. The assessment metrics are detailed in Appendix C.
@@
Most existing assessment metrics (e.g., AUROC, AUPRC, ECE, etc) are rooted in classification and require binary labels (i.e., $A\in\{\textit{True}\text{ or }\textit{False}\}$). Consequently, an ad hoc threshold $\tau\in\mathbb{R}$ is often introduced to map continuous correctness values to binary labels, i.e., $\bar{A}_{\tau}(\boldsymbol{x};\widehat{\boldsymbol{y}}):=\mathds{1}[A(\boldsymbol{x};\widehat{\boldsymbol{y}})\geq\tau]$ Lin et al. (2023); Kuhn et al. (2023). Thus, the response is viewed as correct if the correctness value $A(\boldsymbol{x};\widehat{\boldsymbol{y}})$ is at least $\tau$, and incorrect otherwise.
@@
The second limitation of existing assessments is rooted in the diverse output ranges of the uncertainty or confidence measures. As shown in Fig. 3 [bottom], the output ranges of different uncertainty measures vary significantly. For example, the values of $U_{\rm SE}$ can be higher than $100$ while the values of $U_{\rm Ecc}$ and $U_{\rm Deg}$ are small by definition. This diversity of output ranges prevents the direct use of calibration-based metrics such as ECE, which takes variables with inputs in $[0,1]$.
@@ 4.1 Rank-Calibration & RCE
Since $U$ is a random variable depending on $(\boldsymbol{x};\widehat{\boldsymbol{y}})$, $\mathrm{reg}(U)$ is also random. If $\mathrm{reg}(\cdot)$ is monotonically decreasing, then $U\leq u$ implies $\mathrm{reg}(U)\geq\mathrm{reg}(u)$. Thus, for any value $u$ in the range of $U$,
$\mathbb{P}(U\leq u)=\mathbb{P}(\mathrm{reg}(U)\geq\mathrm{reg}(u)).$ (7)
@@
We say that an uncertainty measure $U$ is rank-calibrated if (7) holds for any $u$ in $U$’s range: on average, a lower uncertainty implies a higher generative quality.
Rank-calibration is related to, yet distinct from, the usual notion of calibration in the classification context (Lichtenstein et al., 1977; Guo et al., 2017). We defer the detailed discussion to Sec. 4.2. We remark that the principle of rank calibration is also discussed in a concurrent work (Zhang et al., 2024). Unlike our work and Zhang et al. (2024), Penha and Hauff (2021) use the terminology “rank” to denote the relevance comparison of candidate responses in the binning of ECE calculation.
@@ 4.2 Comparison with Classical Calibration
Theorem 1 implies that, for a binary correctness function, one can construct a calibrated confidence measure from an uncertainty measure with monotone transformations if and only if the uncertainty measure is rank-calibrated. However, RCE and ECE gauge different quantities: ECE captures the absolute difference between the predicted and true probabilities, while RCE reflects the deviation from a monotonic correspondence between uncertainty and the expected correctness. These two notions are generally not directly comparable.
For example, consider the special case where a continuous-valued confidence measure $C$ is completely uninformative and the regressed correctness $\overline{\mathrm{reg}}:c\mapsto\mathbb{E}[A\mid C=c]$ is a constant for all confidence levels $c$. Then, the RCE defined in (8) reports a large value of $1/2$, reflecting its poor indicativeness. However, the ECE can be large or small depending on the averaged distance between $C$’s output and $\overline{\mathrm{reg}}$. More generally, we find no relation in the results of ECE and RCE through the following result, proved in Appendix D.
@@ 4.3 Empirical RCE & Indication Diagram
To address this, we adopt a piecewise constant regression or binning strategy, as in non-parametric statistics Tsybakov (2009). First, we group the uncertainty values $\{u_{i}\}_{i=1}^{n}$ into $B$ equal-mass intervals, each containing $\lceil n/B\rceil$—or, when needed, $\lfloor n/B\rfloor$—elements. The boundaries of the $b$-th ($1\leq b\leq B$) bin are the $(b-1)/B$-th and $b/B$-th quantiles of $(u_{i})_{i=1}^{n}$. Let $\mathcal{I}_{b}\subseteq\{1,\dots,n\}$ be the set of indices of the datapoints whose uncertainty values fall into the $b$-th bin. The expected correctness level over the $b$-th bin can be estimated as
${\rm crc}_{b}:=\frac{1}{|\mathcal{I}_{b}|}\sum_{i\in\mathcal{I}_{b}}a_{i},$
@@
${\rm uct}_{b}=\frac{1}{|\mathcal{I}_{b}|}\sum_{i\in\mathcal{I}_{b}}u_{i}.$
As ${\rm crc}_{b}$ and ${\rm uct}_{b}$ estimate the per-bin averages of $\mathrm{reg}(U)$ and $U$, for each $b$, we estimate $\mathbb{P}(U\leq u_{i})$ and $\mathbb{P}(\mathrm{reg}(U)\geq\mathrm{reg}(u_{i}))$ for $i\in\mathcal{I}_{b}$ as follows:
^$\widehat{\mathbb{P}}(\mathrm{reg}(U)\geq\mathrm{reg}(u_{i})):=\frac{1}{B-1}\sum_{b^{\prime}\neq b}\mathds{1}[{\rm crc}_{b^{\prime}}\geq{\rm crc}_{b}],\vspace{-1mm}$
^$\widehat{\mathbb{P}}(U\leq u_{i}):=\frac{1}{B-1}\sum_{b^{\prime}\neq b}\mathds{1}[{\rm uct}_{b^{\prime}}\leq{\rm uct}_{b}].$
@@
Dataset Correctness Temperature $U_{\rm Ecc}$ $U_{\rm Deg}$ $U_{\rm EigV}$ $U_{\rm NLL}$ $U_{\rm SE}$ $C_{\rm Verb}$
nq-open bert 0.6 0.199±0.040 0.046±0.008 0.052±0.010 0.101±0.015 0.062±0.010 nan
Table 2: RCE results for Llama-2-chat with various experimental configurations.
@@ 5.1 Experiment Setup
The uncertainty/confidence measures to be assessed are the same as in Sec. 3, (i.e., $U_{\rm NLL}$, $U_{\rm SE}$, $U_{\rm Ecc}$, $U_{\rm Deg}$, $U_{\rm EigV}$, and $C_{\rm Verb}$). We first illustrate that our proposed assessment has broad applicability and granular interpretability. Furthermore, we qualitatively show that uncertainty measures with lower RCE values reliably indicate correctness. Finally, we study robustness by empirically checking the impact of temperature and correctness functions on RCE (Demšar, 2006). More results for different configurations are given in Table 4.
@@ 5.5 Post-hoc Recalibration
Recalibrating uncertainty/confidence measures with poor rank-calibration can be of interest; for ECE, this is sometimes known as Mincer-Zamowitz regression Mincer and Zarnowitz (1969). As discussed in Sec. 4.2, an ECE-calibrated measure is also RCE-calibrated. However, RCE is invariant to monotone transformations, which means that approaches like Platt scaling (Platt, 1999) and isotonic regression (Zadrozny and Elkan, 2002) will not improve rank-calibration. Therefore, we suggest using histogram binning (or, piecewise constant regression), which includes non-monotone transforms (Zadrozny and Elkan, 2001). Table 3 and Fig. 10 and 11 list the RCE results of $U_{\rm SE}$ for GPT-3.5-turbo before and after calibration. We observe the calibrated measure is significantly better rank-calibrated, showing the effectiveness of this strategy. See the more experimental details and results in Appendix F.2.
@@ 5.6 Robustness Analysis
We conduct ablation studies to analyze the robustness of our assessment to key hyperparameters, including temperatures, correctness scores, and sample sizes. We further propose a method to make robust comparisons between uncertainty measures via the Critical Difference (CD) Diagram  (Demšar, 2006). Detailed information and results are in Appendix F.4.
Dataset Correctness Temperature $U_{\rm SE}$ $U_{\rm SE,cal}$
meadow bert 1.0 0.177±0.027 0.083±0.016
meteor 1.0 0.132±0.018 0.066±0.015
rougeL 1.0 0.113±0.022 0.063±0.014
rouge1 1.0 0.113±0.018 0.061±0.012
nq-open bert 1.0 0.050±0.007 0.026±0.007
meteor 1.0 0.060±0.009 0.033±0.011
rougeL 1.0 0.052±0.008 0.030±0.010
rouge1 1.0 0.051±0.008 0.029±0.010
squad bert 1.0 0.113±0.013 0.050±0.013
meteor 1.0 0.086±0.014 0.046±0.010
rougeL 1.0 0.100±0.011 0.037±0.008
rouge1 1.0 0.103±0.011 0.039±0.007
triviaqa bert 0.5 0.052±0.009 0.030±0.010
1.0 0.052±0.012 0.027±0.008
1.5 0.081±0.009 0.029±0.007
meteor 0.5 0.234±0.019 0.058±0.015
1.0 0.209±0.012 0.047±0.014
1.5 0.176±0.015 0.036±0.012
rougeL 0.5 0.050±0.008 0.028±0.007
1.0 0.059±0.009 0.026±0.007
1.5 0.104±0.007 0.028±0.006
rouge1 0.5 0.050±0.008 0.028±0.006
1.0 0.060±0.009 0.027±0.006
1.5 0.105±0.008 0.028±0.008
@@ 6 Conclusion
Appendix A Additional Related Work
@@
The quantification of uncertainties in model outputs in supervised learning has a long history (e.g., Lichtenstein et al., 1977, etc). Overparametrized models such as neural networks pose unique challenges to estimate uncertainty and improve model calibration(Guo et al., 2017; Papadopoulos et al., 2001; Riquelme et al., 2018). Various approaches have been introduced to mimic Bayesian inference (Gal and Ghahramani, 2016), to utilize simple deep ensembles (Lakshminarayanan et al., 2017; Jain et al., 2020), and to identify training samples that are out-of-distribution (Liang et al., 2018; Papernot and McDaniel, 2018). Nonetheless, it is not clear how to adapt these strategies to language modeling, where the output can be text with complex structure.
@@
To gauge the uncertainty level associated with the outputs of LMs, Kuhn et al. (2023) introduces the concept of semantic entropy, which integrates linguistic consistencies stemming from shared meanings. In a similar vein, Kadavath et al. (2022); Lin et al. (2022); Xiong et al. (2024) encourage LMs to analyze their own responses and come up with a “probability” that a response is correct. In related work, Manakul et al. (2023) uses sampling to identify instances of fabricated information. Recently, Tian et al. (2023) explore methods for deriving confidence measures for reinforcement-learning-trained LMs. Lin et al. (2023) draw a distinction between estimating uncertainty and confidence for LMs. Similarly, Chen and Mueller (2023) introduce a method for detecting bad and speculative responses from a pre-trained LM with a confidence score. Tanneru et al. (2023) propose two novel measures to quantify the uncertainty of LM-generated explanations. Although considerable research focuses on developing uncertainty and confidence measures for LMs, the evaluation of their effectiveness is less studied.